OpenStudy (anonymous):

find the value of y for a given value of x. y=x+3, x=13

jimthompson5910 (jim_thompson5910):

x = 13, so replace x with 13 to get y = x+3 y = 13+3 y = ??

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks much

jimthompson5910 (jim_thompson5910):

you're welcome